# CropVLM: Learning to Zoom for Fine-Grained Vision-Language Perception

### v1

@@ 4.2.1 Evaluating CropVLM Paired with SmolVLM
Post-SFT Performance.
Post-GRPO Performance.
Applying GRPO results in substantial improvements across all resolutions and reward types. The 512×512 model has the largest relative gains, where cropping helps compensate for the model’s lower perceptual capacity and visual fragmentation. Log-likelihood rewards generally outperform accuracy-based ones, likely due to the fact that they can provide more nuanced feedback during training. This option virtually eliminates examples where the reward is the same within a group, which in turn leads to more examples effectively contributing to weight updates.
Key Insights.
4.2.2 Evaluating CropVLM with Other VLMs
@@
Table 5: Comparison of our CropVLM approach against other similar approaches in the literature. The CropVLM model present in this table receives as input an image at 2048x2048 pixels and was trained using a log-likelihood reward. Both LLaVA 1.5 and UV-CoT operate at 336x336 pixels of input resolution and Qwen 2.5 VL operates at 448x448 pixels of input resolution. The Average column shows the average performance across datasets.
@@ 4.2.3 Comparing the CropVLM Model Against Alternative Cropping Strategies
We compared our CropVLM strategy against two alternative methods: ViCrop [zhang2025mllms] and UV-CoT [zhao2025unsupervised]. ViCrop is a training-free approach that uses explainability techniques, based on gradients and attention weights, to identify regions of interest within an image, given a query. We evaluate ViCrop using the two best-performing techniques, as reported in the original paper - relative attention (rel-attn) and gradient-weighted attention (grad-attn) - implementing them with the officially released code and recommended hyperparameters, with two base VLMs respectively employing LLaVA 1.5 7B at 336x336 pixels, and Qwen 2.5 at 448x448 pixels resolution, to both generate bounding boxes and answer questions. UV-CoT, on the other hand, is a DPO based training method to learn both a relevant crop and the correct response to a request given an image. We use the released model [^3] https://huggingface.co/kesenZhaoNTU/UV-CoT to produce relevant crops, and pair it with the same answering models as ViCrop. In turn, CropVLM processes inputs at 2048×2048 pixels resolution, while the answering models receive images at the same resolution as their alternative counterparts. Note that VisRL [chen2025visrl] was not available at the time of writing and therefore was not included in our evaluation. Results are presented in Table 5.
@@ 4.3 Execution Overhead
In Table 6, we report the average execution time overhead, and the peak GPU memory usage of CropVLM, comparing these results to those of ViCrop and UV-CoT, as measured on the same 100 random samples from the TextVQA dataset. All measurements were taken on a machine equipped with an NVIDIA RTX A6000 GPU and an Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz. CropVLM demonstrates significantly lower GPU memory usage, largely due to the smaller model size and its compatibility with FlashAttention [dao2023flashattention]. In contrast, ViCrop requires access to the full attention scores, making it incompatible with FlashAttention and resulting in higher memory consumption. On the other hand, CropVLM is slower in generating bounding boxes because it relies on text generation to produce coordinate values, while ViCrop does not. The underlying model (i.e., SmolVLM 256M Instruct) has a highly restricted numeric output vocabulary, limited to the digits 0 through 9. As a result, generating numbers between 0 and 100 requires emitting multiple tokens per number, which increases decoding time. UV-CoT, while also compatible with FlashAttention, is substantially larger and also requires generating multiple tokens for a single bounding box, which makes it slower and more memory-demanding than CropVLM, even at lower resolutions. Future work can perhaps explore alternative approaches to the design of the cropping network.
@@ Appendix B Evaluating the Generated Crops Against Human-annotated Bounding Boxes
  • Intersection over Union (IoU): Measures the overlap between $B_{p}$ and $B_{gt}$.

### v2

@@ 4.2.2 Evaluating CropVLM with Other VLMs
Table 5: Comparison of our CropVLM approach against other similar approaches in the literature. The CropVLM model present in this table receives as input an image at 2048x2048 pixels and was trained using a log-likelihood reward. Both LLaVA 1.5 and UV-CoT operate at 336x336 pixels of input resolution, and Qwen 2.5 VL operates at 448x448 pixels of input resolution. The Average column shows the average performance across datasets.
@@ 4.2.3 Comparing the CropVLM Model Against Alternative Cropping Strategies
We compared our CropVLM strategy against two alternative methods: ViCrop [29] and UV-CoT [31]. ViCrop is a training-free approach that uses explainability techniques, based on gradients and attention weights, to identify regions of interest within an image, given a query. We evaluate ViCrop using the two best-performing techniques, as reported in the original paper - relative attention (rel-attn) and gradient-weighted attention (grad-attn) - implementing them with the officially released code and recommended hyperparameters, with two base VLMs respectively employing LLaVA 1.5 7B at 336x336 pixels, and Qwen 2.5 at 448x448 pixels resolution, to both generate bounding boxes and answer questions. UV-CoT, on the other hand, is a DPO based training method to learn both a relevant crop and the correct response to a request given an image. We use the released model [^3] https://huggingface.co/kesenZhaoNTU/UV-CoT to produce relevant crops, and pair it with the same answering models as ViCrop. In turn, CropVLM processes inputs at 2048×2048 pixels resolution, while the answering models receive images at the same resolution as their alternative counterparts. Note that VisRL [5] was not available at the time of writing and therefore was not included in our evaluation. Results are presented in Table 5.
@@ 4.3 Execution Overhead
In Table 6, we report the average execution time overhead, and the peak GPU memory usage of CropVLM, comparing these results to those of ViCrop and UV-CoT, as measured on the same 100 random samples from the TextVQA dataset. All measurements were taken on a machine equipped with an NVIDIA RTX A6000 GPU and an Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz. CropVLM demonstrates significantly lower GPU memory usage, largely due to the smaller model size and its compatibility with FlashAttention [7]. In contrast, ViCrop requires access to the full attention scores, making it incompatible with FlashAttention and resulting in higher memory consumption. On the other hand, CropVLM is slower in generating bounding boxes because it relies on text generation to produce coordinate values, while ViCrop does not. The underlying model (i.e., SmolVLM 256M Instruct) has a highly restricted numeric output vocabulary, limited to the digits 0 through 9. As a result, generating numbers between 0 and 100 requires emitting multiple tokens per number, which increases decoding time. UV-CoT, while also compatible with FlashAttention, is substantially larger and also requires generating multiple tokens for a single bounding box, which makes it slower and more memory-demanding than CropVLM, even at lower resolutions. Future work can perhaps explore alternative approaches to the design of the cropping network.
@@ 5 Conclusions
Acknowledgements
This research was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (i.e., the Center For Responsible AI), and also by the Fundação para a Ciência e a Tecnologia, I.P. (FCT) under the projects with references UID/50021/2025 (DOI: https://doi.org/10.54499/UID/50021/2025) and UID/PRR/50021/2025 (DOI: https://doi.org/10.54499/UID/PRR/50021/2025).
@@ Appendix B Evaluating the Generated Crops Against Human-annotated Bounding Boxes
  • Intersection over Union (IoU): Measures the relative area of overlap between $B_{p}$ and $B_{gt}$.